# Name That Part: 3D Part Segmentation and Naming

### v1

@@ 4.2. Comparison on Named 3D Part Segmentation
Baselines. While no prior work addresses named 3D part segmentation end-to-end, we compare against two recent baselines: PartField (liu2025partfield) and Find3D (ma2024find). PartField achieves state-of-the-art class-agnostic segmentation but cannot name parts. It clusters a learned feature field (requiring ground-truth part count K). Find3D matches per-point features to SigLIP embeddings of provided part queries, needing the list of ground-truth part names as input.
@@
PartField+MPNet (Our Strong Baseline; w/o partlets). We extend PartField with semantic alignment: a linear head maps PartField features to a shared space with MPNet text embeddings, trained via InfoNCE loss. However, it still requires predicting K via an auxiliary classifier, making it brittle to over/under-segmentation from K prediction errors. Further architecture details are provided in Section˜A.1.
@@
Figure 6. Proposed Metrics Correlation Analysis. Correlation analysis between our proposed label-aware mIoU metrics (strict/relaxed) and class-agnostic mIoU, computed on segmentation results from our ALIGN-Parts model. The strict label-aware metric (left) shows moderate agreement with class-agnostic mIoU (Pearson $r=0.739$, Spearman $\rho=0.730$, $N=206$), while the relaxed variant (right) demonstrates near-perfect correlation (Pearson $r=0.978$, Spearman $\rho=0.974$, $N=206$). These findings indicate that our model achieves strong semantic and quantitative consistency, further supporting the use of the relaxed metric as a robust evaluation protocol for semantic 3D part segmentation.
@@
Figure 7. Label Mismatches and Metric Robustness. Distribution of test objects according to the number of label mismatches (x-axis) and the resulting increase in mIoU from strict to relaxed label-aware matching (y-axis), colored by class-agnostic mIoU. Most objects with low mismatches exhibit small or moderate increases in mIoU, while objects with higher mismatch counts still do not show large outliers in metric difference, supporting the robustness of our evaluation protocol. The absence of extreme discrepancies suggests that the relaxed metric yields a stable and meaningful improvement, even for challenging cases, confirming its reliability for assessing semantic part segmentation on our test set.
@@
Metrics. We evaluate our method using three complementary metrics that progressively incorporate semantic label correctness. (1) Class-agnostic mIoU: Following prior work (liu2025partfield; yang2024sampart3d), for each ground-truth part, we compute the maximum IoU across all predicted segments and average these values, ignoring semantic labels entirely - this captures pure geometric segmentation quality. (2) Label-Aware mIoU (strict, LA-mIoU): For each ground-truth part, we identify the predicted segment with the highest geometric overlap (as in class-agnostic mIoU), then assign credit only if its semantic label exactly matches the ground truth; otherwise, the part contributes 0.0 - this measures joint geometry-semantic accuracy with strict label matching. (3) Relaxed Label-Aware mIoU (rLA-mIoU): Identical segment selection to strict LA-mIoU, but instead of binary label matching, we weight the IoU by the cosine similarity between MPNet text embeddings of predicted and ground-truth labels, giving partial credit to semantically related predictions (e.g., "screen" vs. "monitor") - this captures semantic near-misses that strict matching penalizes. By construction, we have class-agnostic mIoU $\geq$ rLA-mIoU $\geq$ LA-mIoU (strict), where equality holds only when all predictions have perfect label agreement. The gap between class-agnostic and label-aware metrics reveals semantic prediction errors, while the gap between strict and relaxed variants quantifies label confusion on semantically similar parts. Correlation plots (Fig 6 and 7) empirically validate that relaxed scoring recovers significant semantic credit on near-miss predictions.
4.3. Results
@@
Table 2. Evaluation of ALIGN-Parts and related baselines on our test set. In addition to the usual mIoU metric for evaluating class-agnostic part segmentation, we introduce 2 metrics more suited for our named part-segmentation task, namely, label-aware mIoU (LA-mIoU) and relaxed label-aware mIoU (rLA-mIoU).
@@
In Figures 5 and 8 and Table 2, we show that ALIGN-Parts comprehensively outperforms all baselines on both class-agnostic segmentation (mIoU) and named part segmentation (LA-mIoU). On average mIoU, we outperform PartField by 15.8%, whereas on LA-mIoU and rLA-mIoU, we improve over PartField+MPNet by 58.8% and 43.8%, respectively, while being $100\times$ faster because we do not require running a K-means clustering algorithm.
Qualitatively, our baselines show several weaknesses. PartField often fragments instances of the same part into multiple segments after clustering, contradicting human labeling conventions in 3DCoMPaT++ and PartNet. This likely stems from its use of SAM (kirillov2023segany) to extract unlabeled parts from 2D renderings. In contrast, ALIGN-Parts, trained on human annotations, correctly groups instances: e.g., all four bed posts of a double bed and all three wheels of an airplane are identified as single parts (Figure 8).
Find3D learns per-point semantic vectors without considering shape geometry, resulting in noisy and overlapping segmentations. PartField+MPNet’s reliance on predicted cluster counts leads to under-segmentation, as fine parts such as the refrigerator handle and laptop screen frame are missed (Figure 8). Both Find3D and PartField fail to segment even simple structures, such as fan blades, whereas ALIGN-Parts accurately segments and annotates fine-grained parts in complex shapes.
Runtime Comparison. In terms of runtime, our method compares favorably against all baselines. Find3D runs in $\approx 0.25s$, whereas both PartField and PartField+MPNet require $\approx 4s$, where the majority of the runtime is consumed by K-means clustering. ALIGN-Parts, being a one-shot feedforward method, needs only $\approx 0.05s$ (barring feature pre-processing) to produce labelled 3D parts.
Table 3. Ablation study on 3DCoMPaT, Find3D, and PartNet. We report mIoU, label-aware mIoU (LA-mIoU), and relaxed label-aware mIoU (rLA-mIoU).
@@
Fine-Part Localization Despite using only 10k sampled points, ALIGN-Parts segments fine parts like the screw of scissors (Fig 8) - structures that PartField, trained with 100k points, cannot localize.
Figure 8. ALIGN-Parts correctly segments the tiny screw despite training with sparser points (10k vs. 100k).
Figure 9. Ablation (Qualitative results). Results improve from left to right as components are added sequentially, with ground truth in the final column. The first five columns result in significant misalignments and segmentation leakage. Neither geometric-only features nor naive DINO concatenation improves performance. Major gains arise from Partlets (see sixth and seventh columns) and the coverage loss, which refine fine details and are consistent with quantitative metrics in Table 3.
4.4. Ablations
Table 3 and Figure 9 evaluate ALIGN-Parts’s design choices, showing quantitative and qualitative improvements from each component.
Baseline Comparisons. Using only PartField geometric features, our base model ($\mathcal{L}_{\text{mask}},\mathcal{L}_{\text{part}},\mathcal{L}_{\text{global}}$) performs slightly better than naive concatenation of PartField and DINOv2 features, indicating that simple multi-modal fusion can be counterproductive. A variant using raw geometry without PartField shows no improvement, confirming that learned geometric features are essential.
Progressive Component Addition. Adding DINOv2 appearance features yields modest gains. Incorporating the InfoNCE text alignment loss ($\mathcal{L}_{\text{txt}}$) significantly improves both LA-mIoU and rLA-mIoU; this variant is the only one, besides our full model, that correctly segments fine-grained parts, such as plants in vases or shelves in storage furniture. Finally, adding auxiliary regularizers ($\mathcal{L}_{\text{cov}},\mathcal{L}_{\text{overlap}}$) yields the complete ALIGN-Parts model, which achieves peak performance both qualitatively and quantitatively.
4.5. Inference-time Ablations
Table 4. Evaluation of different inference modes of ALIGN-Parts on our test set, using mean IoU (mIoU), label-aware mIoU (LA-mIoU), and relaxed label-aware mIoU (rLA-mIoU). Providing additional ground-truth part count information only slightly improves the model’s performance on rLA-mIoU showing that ALIGN-Parts often estimates accurate part cardinality based on just the input geometric and appearance features of a 3D shape.
@@ 4.5. Inference-time Ablations
Figure 10. Inference-time Ablations. We assess two additional inference modes in our ALIGN-Parts model. The Clustering+Part number mode disregards text vocabularies and applies k-means to the fused geometric and appearance features, generating class-agnostic clusters and probing how well raw feature representations alone support meaningful part segmentation. The +Part number mode uses the true part count for each shape, activating exactly $M$ partlets with the highest saliency scores - computed by combining partness confidence and average mask coverage - and assigning every point to its best matching mask; this tests whether supplying ground-truth cardinality adds value compared to dynamic, data-driven part discovery. Both setups are compared to the default dynamic activation (No Part number) and the ground truth, highlighting that ALIGN-Parts robustly estimates part cardinality and produces accurate segmentations even without explicit part label or count supervision.
@@
We evaluate two additional inference modes of our ALIGN-Parts model, extending beyond the primary dynamic part activation approach to better understand the contributions of part cardinality and label information in our segmentation pipeline. The first alternative mode, which we term the clustering+part number setting, completely forgoes the use of any part vocabularies or text labels during inference. Instead, it relies solely on the fused geometric and appearance features output by the model, upon which we run k-means clustering to produce purely class-agnostic instance clusters. This setup rigorously probes the ability of the learned feature embeddings, untethered to semantic labels, to support coherent part decompositions across diverse objects, essentially isolating the impact of visual and geometric cues alone. The second mode, called +Part number, examines whether providing the model with the exact ground-truth part count for each input shape improves segmentation quality compared to the default setting, where the model dynamically infers the number of parts to activate. After producing all candidate partlet masks and calculating their partness scores, this mode ranks the partlets by a saliency score, which is a composite measure combining the confidence that a partlet corresponds to an actual part (i.e., partness) and the average mask coverage over the point cloud (mean mask probability mass over points). From this ranking, the top $M$ partlets are retained, where $M$ is the true number of parts for the target shape, and every point in the shape is assigned the best matching mask among these selected partlets to yield a hard $K$-way partition. These inference ablation modes and their qualitative and quantitative outcomes are detailed and visualized in Fig 10 and Tab 4, demonstrating that ALIGN-Parts is able to robustly estimate accurate part cardinality and segmentation even without explicit part label or count guidance, and that the fused multimodal features alone provide meaningful cues towards coherent part delineation. This analysis not only highlights the flexibility and robustness of ALIGN-Parts at inference time, but also emphasizes the benefits of its design choices in learning and leveraging rich feature representations supporting both semantic and instance-aware part segmentation.
Figure 11. 3D Shape Co-Segmentation Analysis. Left: clustering-based co-segmentation. Prior methods such as PartField perform dependent clustering by first segmenting a source shape via feature clustering and then using the resulting cluster means to initialize K-means on a target shape, implicitly enforcing part correspondence; this can break when the target has a different part count or geometry, causing errors such as the red backrest region bleeding into the seat on the target chair. Using the same dependent co-segmentation strategy with our fused BiCo features (BiCo Feature Dependent Clustering) yields improved transfers on moderately similar targets (Target A), but performance degrades on more challenging targets with greater variation in part structure (Target B). As an alternative, we apply independent clustering to Target C, where the target is segmented with source initialization and clusters are matched post hoc by comparing source and target cluster centers, which proves more reliable for difficult co-segmentation cases. Middle and right: feedforward ALIGN-Parts. In contrast to all clustering-based variants, the proposed feedforward ALIGN-Parts model (middle) directly predicts part segmentation and names, achieving robust results across shapes with differing part counts and topologies, and eliminating any dependence on source shapes or explicit co-segmentation.
4.6. 3D Shape Co-Segmentation and Part Label Transfer
Fig.˜11 shows results and analysis of 3D Shape Co-Segmentation using ALIGN-Parts (and the BiCo features) as compared to PartField. A classical approach to 3D part segmentation operates in a co-segmentation setting, where multiple shapes from the same category are jointly analyzed to establish consistent part correspondence. Prior methods, including PartField, employ what we call dependent clustering: they first segment a source shape using feature clustering, then initialize k-means clustering on a target shape using the source cluster centroids, implicitly enforcing part correspondence. While this strategy can succeed on geometrically similar shapes, it proves fragile when target shapes exhibit different part counts or topologies. For example, in Figure 11, dependent clustering on a moderately similar target chair (Target A) produces reasonable results, but fails dramatically on targets with substantial part variation (Target B), causing geometric boundaries to blur (e.g., the backrest merging incorrectly with the seat).
In contrast to both clustering-based paradigms, our proposed ALIGN-Parts adopts a fully feedforward, discriminative approach that predicts part segmentation masks and semantic labels jointly, without requiring source shape initialization or explicit co-segmentation. This design eliminates brittleness to part count variation and geometric mismatches, enabling robust generalization across shapes with diverse part structures and semantics. As demonstrated in Figure 11 (middle and right panels), ALIGN-Parts consistently produces accurate, semantically grounded part segmentations regardless of target shape complexity.
4.7. TexParts Dataset
@@
Figure 12. TexParts Dataset. We demonstrate human-in-the-loop annotation of Texverse (zhang2025texverse) using ALIGN-Parts, enabling scalable dense 3D part segmentation.
@@
A central aim of our approach is to enable the construction of a high-quality 3D part annotation dataset with minimal manual intervention, ensuring both unified and comprehensive part labeling at scale. For this purpose, we select the TexVerse dataset as our unannotated 3D source corpus, leveraging its exceptional quality, high-resolution textures, and extensive diversity of 3D assets (zhang2025texverse). TexVerse consists of over 850,000 unique 3D models with physically based rendering (PBR) materials and rich metadata, making it an ideal foundation for large-scale part segmentation.
@@
During annotation, annotators use several aids: a part name prompting tool for searching or extending the active part vocabulary, and (optionally) the ability to reference unlabeled geometric part masks generated by PartField. Our annotation process is explicitly bilevel; Phase One focuses on validating and making minor edits to ALIGN-Parts predictions, while Phase Two addresses new or missing parts that require more substantive manual annotation. By the time of submitting this work, the first phase has covered approximately 8,000 objects, comprising around 14,000 unique part categories. Examples from the current dataset are shown in Figure 12.
@@
5. Limitations and Future Work

### v2

@@ 4.2 Comparison for Named 3D Part Segmentation
While no prior work addresses named 3D part segmentation end-to-end, we compare against two recent baselines: PartField [23] and Find3D [26]. PartField achieves state-of-the-art class-agnostic segmentation but cannot name parts. It clusters a learned feature field (requiring ground-truth part count K). Find3D matches per-point features to SigLIP embeddings of provided part queries, needing the list of ground-truth part names as input.
PartField+MPNet (our baseline w/o partlets).
@@
We extend PartField with semantic alignment: a linear head maps PartField features to a shared space with MPNet text embeddings, trained via InfoNCE loss. However, it still requires predicting K via an auxiliary classifier, making it brittle to over- or under-segmentation errors in K prediction. Further architecture details are provided in Sec.˜B.1.
@@
Figure 5: Proposed Metrics Correlation Analysis. Correlation analysis between our proposed label-aware mIoU metrics (strict/relaxed) and class-agnostic mIoU, computed on segmentation results from our ALIGN-Parts model. The strict label-aware metric (left) shows moderate agreement with class-agnostic mIoU (Pearson $r=0.739$, Spearman $\rho=0.730$, $N=206$), while the relaxed variant (right) demonstrates near-perfect correlation (Pearson $r=0.978$, Spearman $\rho=0.974$, $N=206$). These findings indicate that our model achieves strong semantic and quantitative consistency, further supporting the use of the relaxed metric as a robust evaluation protocol for semantic 3D part segmentation.
@@
Figure 6: Label Mismatches and Metric Robustness. Distribution of test objects according to the number of label mismatches (x-axis) and the resulting increase in mIoU from strict to relaxed label-aware matching (y-axis), colored by class-agnostic mIoU. Most objects with low mismatches exhibit small or moderate increases in mIoU, while objects with higher mismatch counts still do not show large outliers in metric difference, supporting the robustness of our evaluation protocol. The absence of extreme discrepancies suggests that the relaxed metric yields a stable and meaningful improvement, even for challenging cases, confirming its reliability for assessing semantic part segmentation on our test set.
Metrics.
@@
We evaluate our method using three complementary metrics that progressively incorporate semantic label correctness. (1) Class-agnostic mIoU: Following prior work [23, 42], for each ground-truth part, we compute the maximum IoU across all predicted segments and average these values, ignoring semantic labels entirely - this captures pure geometric segmentation quality. (2) Label-Aware mIoU (strict, LA-mIoU): For each ground-truth part, we identify the predicted segment with the highest geometric overlap (as in class-agnostic mIoU), then assign credit only if its semantic label exactly matches the ground truth; otherwise, the part contributes 0.0 - this measures joint geometry-semantic accuracy with strict label matching. (3) Relaxed Label-Aware mIoU (rLA-mIoU): Identical segment selection to strict LA-mIoU, but instead of binary label matching, we weight the IoU by the cosine similarity between MPNet text embeddings of predicted and ground-truth labels, giving partial credit to semantically related predictions (e.g., "screen" vs. "monitor") - this captures semantic near-misses that strict matching penalizes. By construction, we have class-agnostic mIoU $\geq$ rLA-mIoU $\geq$ LA-mIoU (strict), where equality holds only when all predictions have perfect label agreement. The gap between class-agnostic and label-aware metrics reveals semantic prediction errors, while the gap between strict and relaxed variants quantifies label confusion on semantically similar parts. Correlation plots (Fig.˜5 and Fig.˜6) empirically validate that relaxed scoring recovers significant semantic credit on near-miss predictions.
4.3 Results
@@
Table 2: Evaluation of ALIGN-Parts and related baselines on our test set. In addition to the usual mIoU metric for evaluating class-agnostic part segmentation, we introduce 2 metrics more suited for our named part-segmentation task, namely, label-aware mIoU (LA-mIoU) and relaxed label-aware mIoU (rLA-mIoU).
@@
In Fig.˜4 and Fig.˜7 and Tab.˜2, we show that ALIGN-Parts comprehensively outperforms all baselines on both class-agnostic segmentation (mIoU) and named part segmentation (LA-mIoU). On average mIoU, we outperform PartField by 15.8%, whereas on LA-mIoU and rLA-mIoU, we improve over PartField+MPNet by 58.8% and 43.8%, respectively, while being $100\times$ faster because we do not require running a K-means clustering algorithm.
Qualitatively, our baselines show several weaknesses. PartField often fragments instances of the same part into multiple segments after clustering, contradicting human labeling conventions in 3DCoMPaT++ and PartNet. This likely stems from its use of SAM [18] to extract unlabeled parts from 2D renderings. In contrast, ALIGN-Parts, trained on human annotations, correctly groups instances: e.g., all four bed posts of a double bed and all three wheels of an airplane are identified as semantically single parts (Fig.˜7).
Find3D learns per-point semantic vectors without considering shape geometry, resulting in noisy and overlapping segmentations. PartField+MPNet’s reliance on predicted cluster counts leads to under-segmentation, as fine parts such as the refrigerator handle and laptop screen frame are missed (Fig.˜7). Both Find3D and PartField often fail to segment relatively simple structures, such as fan blades, whereas ALIGN-Parts accurately segments and annotates fine-grained parts in complex shapes.
Runtime Comparison.
In terms of runtime, our method compares favorably against all baselines. Find3D runs in $\approx 0.25s$, whereas both PartField and PartField+MPNet require $\approx 4s$, where the majority of the runtime is consumed by K-means clustering. ALIGN-Parts, being a one-shot feedforward method, needs only $\approx 0.05s$ (barring feature pre-processing) to produce labelled 3D parts.
Table 3: Ablation study on 3DCoMPaT, Find3D, and PartNet. We report mIoU, label-aware mIoU (LA-mIoU), and relaxed label-aware mIoU (rLA-mIoU).
@@
Fine-Part Localization
Despite using only 10k sampled points, ALIGN-Parts segments fine parts, such as the screw of scissors (Fig.˜7) - structures that PartField, trained with 100k points, cannot localize.
Figure 7: ALIGN-Parts correctly segments the tiny screw despite training with sparser points (10k vs. 100k).
Figure 8: Ablation (Qualitative results). Results improve from left to right as components are added sequentially, with ground truth in the final column. The first five columns result in significant misalignments and segmentation leakage. Neither geometric-only features nor naive DINO concatenation improves performance. Major gains arise from Partlets (see sixth and seventh columns) and the coverage loss, which refine fine details and are consistent with quantitative metrics in Tab.˜3.
4.4 Ablations
Tab.˜3 and Fig.˜8 evaluate ALIGN-Parts’s design choices, showing quantitative and qualitative improvements from each component.
Baseline Comparisons.
Using only PartField geometric features, our base model ($\mathcal{L}_{\text{mask}},\mathcal{L}_{\text{part}},\mathcal{L}_{\text{global}}$) performs slightly better than naive concatenation of PartField and DINOv2 features, indicating that simple multi-modal fusion can be counterproductive. A variant using raw geometry without PartField shows no improvement, confirming that learned geometric features are essential.
Progressive Component Addition.
Adding DINOv2 appearance features yields modest gains. Incorporating the InfoNCE text alignment loss ($\mathcal{L}_{\text{txt}}$) significantly improves both LA-mIoU and rLA-mIoU; this variant is the only one, besides our full model, that correctly segments fine-grained parts, such as plants in vases or shelves in storage furniture. Finally, adding auxiliary regularizers ($\mathcal{L}_{\text{cov}},\mathcal{L}_{\text{overlap}}$) yields the complete ALIGN-Parts model, which achieves peak performance both qualitatively and quantitatively.
4.5 Inference-time Ablations
Table 4: Evaluation of different inference modes of ALIGN-Parts on our test set, using mean IoU (mIoU), label-aware mIoU (LA-mIoU), and relaxed label-aware mIoU (rLA-mIoU). Providing additional ground-truth part count information only slightly improves the model’s performance on rLA-mIoU showing that ALIGN-Parts often estimates accurate part cardinality based on just the input geometric and appearance features of a 3D shape.
@@ 4.5 Inference-time Ablations
Figure 9: Inference-time Ablations. We assess two additional inference modes in our ALIGN-Parts model. The Clustering+Part number mode disregards text vocabularies and applies k-means to the fused geometric and appearance features, generating class-agnostic clusters and probing how well raw feature representations alone support meaningful part segmentation. The +Part number mode uses the true part count for each shape, activating exactly $M$ partlets with the highest saliency scores - computed by combining partness confidence and average mask coverage - and assigning every point to its best matching mask; this tests whether supplying ground-truth cardinality adds value compared to dynamic, data-driven part discovery. Both setups are compared to the default dynamic activation (No Part number) and the ground truth, highlighting that ALIGN-Parts robustly estimates part cardinality and produces accurate segmentations even without explicit part label or count supervision.
@@
We evaluate two additional inference modes of our ALIGN-Parts model, extending beyond the primary dynamic part activation approach to better understand the contributions of part cardinality and label information in our segmentation pipeline. The first alternative mode, which we term the clustering+part number setting, completely forgoes the use of any part vocabularies or text labels during inference. Instead, it relies solely on the fused geometric and appearance features output by the model, upon which we run k-means clustering to produce purely class-agnostic instance clusters. This setup rigorously probes the ability of the learned feature embeddings, untethered to semantic labels, to support coherent part decompositions across diverse objects, essentially isolating the impact of visual and geometric cues alone. The second mode, called +Part number, examines whether providing the model with the exact ground-truth part count for each input shape improves segmentation quality compared to the default setting, where the model dynamically infers the number of parts to activate. After producing all candidate partlet masks and calculating their partness scores, this mode ranks the partlets by a saliency score, which is a composite measure combining the confidence that a partlet corresponds to an actual part (i.e., partness) and the average mask coverage over the point cloud (mean mask probability mass over points). From this ranking, the top $M$ partlets are retained, where $M$ is the true number of parts for the target shape, and every point in the shape is assigned the best matching mask among these selected partlets to yield a hard $K$-way partition. These inference ablation modes and their qualitative and quantitative outcomes are detailed and visualized in Fig.˜9 and Tab.˜4, demonstrating that ALIGN-Parts is able to robustly estimate accurate part cardinality and segmentation even without explicit part label or count guidance, and that the fused multimodal features alone provide meaningful cues towards coherent part delineation. This analysis not only highlights the flexibility and robustness of ALIGN-Parts at inference time, but also emphasizes the benefits of its design choices in learning and leveraging rich feature representations supporting both semantic and instance-aware part segmentation.
4.6 TexParts Dataset
@@
Figure 10: TexParts Dataset. We demonstrate human-in-the-loop annotation of Texverse [43] using ALIGN-Parts, enabling scalable dense 3D part segmentation.
@@
A central aim of our approach is to enable the construction of a high-quality 3D part annotation dataset with minimal manual intervention, ensuring both unified and comprehensive part labeling at scale. For this purpose, we select the TexVerse dataset as our unannotated 3D source corpus, leveraging its exceptional quality, high-resolution textures, and extensive diversity of 3D assets [43]. TexVerse consists of over 850,000 unique 3D models with physically based rendering (PBR) materials and rich metadata, making it an ideal foundation for large-scale part segmentation.
@@
During annotation, annotators use several aids: a part name prompting tool for searching or extending the active part vocabulary, and (optionally) the ability to reference unlabeled geometric part masks generated by PartField. Our annotation process is explicitly bilevel; Phase One focuses on validating and making minor edits to ALIGN-Parts predictions, while Phase Two addresses new or missing parts that may require more substantive manual annotation. By the time of submission of this work, the first phase had covered approximately 8,000 objects, comprising around 14,000 unique part categories. Examples from the current dataset are shown in Fig.˜10.
@@
5 Limitations and Future Work